Oxygen initially at 1000psia and 0.4 ft³/lb undergoes an isothermal process to P2 = 200psia. If the process is steady flow with AK = -2 Btu/lb and performs ac cording to the Berthelot equation, determine a, b, -f v dp, and the work W. The mass flow rate is 16 lb/sec.

1
Expert's answer
2021-12-15T11:21:25-0500

"\frac{V_2}{V_1}=\frac{p_1}{p_2}=5,"

"W=A+p_1V_1\ln\frac{p_1}{p_2}=644~\frac{Btu}{lb}."